# RomanSetu: Efficiently unlocking multilingual capabilities of Large Language Models via Romanization

### v2

@@ Abstract
This study addresses the challenge of extending Large Language Models (LLMs) to non-English languages using non-Roman scripts. We propose an approach that utilizes the romanized form of text as an interface for LLMs, hypothesizing that its frequent informal use and shared tokens with English enhance cross-lingual alignment. Our approach involves the continual pretraining of an English LLM like Llama 2 (Touvron et al., 2023) on romanized text of non-English, non-Roman script languages, followed by instruction tuning on romanized data. The results indicate that romanized text not only reduces token fertility by 2x-4x but also matches or outperforms native script representation across various NLU, NLG, and MT tasks. Moreover, the embeddings computed on romanized text exhibit closer alignment with their English translations than those from the native script. Our approach presents a promising direction for leveraging the power of English LLMs in languages traditionally underrepresented in NLP.
@@
[…]
@@ 1 Introduction
Large language models (LLMs) demonstrate remarkable proficiency across a broad spectrum of natural language processing (NLP) tasks, as evidenced by various studies (Liu et al., 2023; Chung et al., 2022; Chowdhery et al., 2022; Wei et al., 2022; Goyal et al., 2022b). They excel not only in tasks for which they were explicitly trained but also in those they were not trained for. This achievement is mainly due to the availability of corpora (Wenzek et al., 2019; Abadji et al., 2021; Suárez et al., 2019) as well as the advancements in LLMs that leverage these datasets for pretraining (Touvron et al., 2023; Workshop et al., 2022; Chowdhery et al., 2022). Despite their proficiency in English, these models typically demonstrate reduced effectiveness when applied to non-English languages, highlighting a significant challenge in extending their benefits to non-English languages.
@@
Figure 1: Performance of the BaseLLM (Llama 2 Base), our continually pretrained model (CPT), and our instruction finetuned model (IFT), in both native (N) and romanized (R) script settings. The average scores for different tasks (left) and languages (right) are compared in each radar chart. For CPT, we show 3-shot results where available, else 1-shot results. For IFT, we use zero-shot results.
@@
The English-heavy LLMs (Touvron et al., 2023; Jiang et al., 2023; Zhang et al., 2022) still have some representation from other languages due to data leakage while creating the pre-training dataset, particularly for languages which use the same script as English i.e., the Roman script. This script sharing enables cross-lingual transfer and bestows some of these LLM capabilities to these languages. For languages using non-Latin scripts, the data representation is very limited to non-existent. Tokenization of text in the languages exhibits high fertility (Ács, 2020) and byte-level representation (Artetxe et al., 2019). Hence, these LLMs perform poorly on most of these languages, and inefficient tokenization also leads to high inference latency as well as inability to process long sequences. This disparity in performance raises a critical question: how can we extend the capabilities of LLMs to the languages written in non-Latin scripts?
@@
A widely explored solution is extension of the tokenizer vocabulary to incorporate new languages and continual pre-training on native language data (Cui et al., 2023; Nguyen et al., 2023; Minixhofer et al., 2021). This approach is computationally demanding, since the models need to be continually pre-trained long enough to effectively learn the new embeddings and align the representations of English and the new language. Furthermore, this approach requires the availability of large volumes of text corpora.
@@
In this work, we explore an alternative approach for more efficient knowledge transfer. Instead of utilizing the native script, we use the romanized form of text as the interface to the LLMs in our approach RomanSetu. [^1] The word setu in Sanskrit means bridge, referring to the Roman script serving as a bridge between English and other languages. The adoption of romanized representation is justified for several reasons. In many languages, romanized text has been frequently used in informal settings and on social media in recent history. This usage creates the potential for the inclusion of some romanized data representation in the pre-training corpus. Additionally, code-mixing with English is a common occurrence, and the romanized form shares tokens with English. This leads us to hypothesize that the romanized form is better aligned with English than the native script, thereby facilitating a more effective transfer from English.
@@
In our approach, we first continually pre-train a English LLM like Llama 2 (Touvron et al., 2023) on romanized text for a language generated via transliteration. Subsequently, we perform instruction tuning on this continually pretrained model. We experiment with multiple languages from two language families and evaluate our approach on multiple NLU, NLG and MT benchmarks.
Our experiments and analysis indicate that:
@@
  • Results across NLU, NLG, and MT tasks shows that multilingual instruction finetuning on romanized data produces results comparable or better than instruction finetuning on native script data, highlighting efficiency and better cross-lingual transfer. Figure 1 summarizes these results of various models on romanized and native inputs.
@@
  • Romanized representation can enable cross-lingual transfer in decoder-only English-heavy LLMs. Previous work has primarily focused on use of romanization in multilingual, encoder-only models.
We will release our code, datasets and model checkpoints upon the paper’s acceptance.
@@ 2 Related Work
In NMT, Amrhein and Sennrich (2020) show that transliteration shows improvements for low-resource languages with different scripts by transferring from related high resource languages that use different scripts. Goyal et al. (2020) show that transliteration helps even when only contact relationship exists between the languages involved. Song et al. (2020) show that transliteration during pre-training stage for NMT also helps cross-lingual transfer.
Transliteration has also been used for cross-lingual transfer in the context of pre-trained language models. Some works (Khemchandani et al., 2021; Dhamecha et al., 2021; Moosa et al., 2022; Purkayastha et al., 2023) employ transliteration to a common script during the pretraining phase to enable cross-lingual transfer. Other studies (Dabre et al., 2022; Muller et al., 2021; Chau and Smith, 2021) adopt transliteration during the fine-tuning. Transliteration could be done to a common non-Latin script (Khemchandani et al., 2021; Dhamecha et al., 2021; Dabre et al., 2022; Doddapaneni et al., 2023) or to the Latin script Muller et al. (2021); Moosa et al. (2022); Purkayastha et al. (2023).
@@
  • Previous work explored cross-lingual transfer using transliteration with multilingual language models. We focus our attention on English LLMs, and try to achieve cross-lingual transfer via romanization using English LLMs. This is a challenging scenario, since the language can be unrelated to English, and very little (if any) native or romanized data in the language under study might be seen during pre-training of the English LLM. At the same time, this is a very practical need since most best performing LLMs are English-heavy Touvron et al. (2023); Jiang et al. (2023) and hence cross-lingual transfer via romanization is an important direction to explore.
@@
To the best of our knowledge, no prior research has investigated leveraging of romanization for mostly-English LLMs in the context of cross-lingual transfer to non-English languages.
@@ 3 Utilizing Romanized Data to make LLMs Multilingual
Our proposed approach aims to enhance the capabilities of mostly-English LLMs for non-English languages by leveraging romanized data. We first continually pretrain the LLM with both romanized and English data to create a base LM that is romanization-aware. Subsequently, we extend the process by instruction tuning the continually pretrained model. The framework of our approach is illustrated in Figure 2, encompassing the stages of data romanization, pre-training, and instruction tuning.
@@ 3.1 Romanization Scheme
A variety of romanization schemes are available, each driven by multiple considerations. One key factor is the resemblance of the romanized representation to the way people typically write romanized text. This is particularly advantageous if the pre-training data includes romanized text, as it might aid in aligning with English. Another important aspect is the fertility achieved by the romanization scheme when considering the original LLM’s tokenizer. A third consideration is whether the transliteration scheme is lossy or lossless. A lossless scheme is preferable when the objective is to convert the output back to the native script. Typically, deterministic transliteration schemes are lossless, whereas natural transliteration schemes tend to be lossy.
In this work, we have focused on Indic languages. Hence, we evaluated two romanization schemes for Indic languages: (a) the extended ITRANS scheme from the IndicNLP library Kunchukuttan (2020), which defines a fixed, reversible mapping between Indic scripts and Roman characters, and (b) the IndicXlit scheme Madhani et al. (2023), which generates romanizations as commonly used by Hindi speakers in informal contexts and is learned from parallel transliteration corpora. These mappings are inherently lossy. The IndicXlit romanization demonstrates lower fertility compared to ITRANS romanization (See Table 1). Preliminary prompting experiments with the Llama 2 model also indicated that IndicXlit outperforms ITRANS in romanized Hindi to English translation. Consequently, we selected IndicXlit as our romanization scheme for this project. However, it is important to note that the transliterations are not reversible. With continued pre-training, ITRANS transliterations might eventually achieve similar task performance to IndicXlit while ensuring script reversibility. This direction might also be relevant for other languages which do not have a transliteration model available, while romanizations mappings are available for almost all scripts. We leave this exploration for future work.
@@ 3.2 Continual Pretraining
While the base model is trained on roman script, most of the data it has been exposed to is in English. Hence, the base model cannot be used as-is for processing romanized inputs, particularly for generation into non-English languages. Hence, the base model needs to be updated. To make the base model romanization-aware, we continue to pre-train the base LLM with romanized document-level data. To prevent any catastrophic forgetting of English capabilities, we also incorporate an equal amount of English data into the pre-training mix. We perform full-finetuning of the models. We do not expand or change the model vocabulary, hence the romanized inputs can take advantage of the token embeddings from the base model while adapting to the new languages.
@@ 3.3 Instruction Finetuning (IFT)
Non-English languages possess very limited finetuning data. A common method to overcome this limitation is to translate English IFT datasets into the languages. We could choose to translate both input and outputs Wei et al. (2023), or translate only the inputs and ask a powerful, proprietary LLM like ChatGPT to generate translations in the native language Li et al. (2023). We chose the former approach since the quality of LLM responses (even for the best commerical models) might not be of good quality Ahuja et al. (2023), while high-quality open-source translation models Costa-jussà et al. (2022); Gala et al. (2023) are available making translation of both inputs and outputs feasible at scale for many languages. The translated IFT datasets are then romanized using IndicXlit.
@@ 4.1 Languages
We experimented with the following Indic languages: Hindi, Marathi, Gujarati, Tamil and Malayalam. These languages span 2 languages families and 4 different scripts and are mid to low-resource languages. Hence these are representative of various multilingual scenarios where romanization can be put to use.
@@ 4.2 Datasets
Instruction Fine-tuning: These languages have very little native instruction tuning data for diverse tasks. Following Wei et al. (2023), we rely on translating high-quality English-supervised instruction-tuning datasets into the langauges under consideration. We use IndicTrans2 (Gala et al., 2023), the state-of-the-art open-source MT model for Indian languages compared to commercial offerings for translation. We sampled examples from various English instruction tuning datasets to ensure a diverse mix of tasks from the Flan collection (Longpre et al., 2023) (65k) and Dolly (Conover et al., 2023) (15k). These are translated into all the languages. The instruction tuning dataset includes a translation subset as well to learn the translation task as well as to drive alignment between English and other languages. We sample 20k high-quality manually translated examples in each direction from the BPCC-Human subset (Gala et al., 2023). Thus, the final IFT dataset has 120k examples per language. The instruction tuning datasets are further romanized using IndicXlit.
Evaluation Data: We evaluate our models on a variety of NLU and NLG tasks. These include native language test sets from the IndicXTREME (Doddapaneni et al., 2023) and Indic NLG Suite (Kumar et al., 2022) benchmarks. For the translation task, we use the FLORES-200 devtest (Goyal et al., 2022a; Costa-jussà et al., 2022) for evaluation. Following recent work (Lai et al., 2023; Üstün et al., 2024; OpenAI et al., 2023), we evaluate the knowledge and reasoning capabilities of the model on the machine-translated versions of the benchmarks such as MMLU (Hendrycks et al., 2021a), ARC (Clark et al., 2018a), BoolQ (Clark et al., 2019a) and CommonSenseQA (Talmor et al., 2019). Note that as the CommonSenseQA benchmark comprises of a blind test set, we report the evaluations on the dev set. The translations were obtained using IndicTrans2 and the romanized inputs are created using IndicXlit.
@@ 4.3 Training and Finetuning Details
The models are fully-finetuned during continual pre-training as well supervised finetuning. Both the continued pre-training and instruction finetuning is multilingual i.e. a single model is trained jointly with data for all languages (including English to preserve English performance). Note that no vocabulary expansion is done. For pre-training, loss is computed for all the tokens in the sequence. During instruction fine-tuning, the loss was computed only for the output tokens. Detailed information about the hyperparameters used for both continual pretraining and instruction fine-tuning can be found in Appendix A. We adapt open-instruct [^2] https://github.com/allenai/open-instruct/tree/main (Wang et al., 2023) for our continual pre-training and finetuning experiments.

### v3

@@ Abstract
This study addresses the challenge of extending Large Language Models (LLMs) to non-English languages that use non-Roman scripts. We propose an approach that utilizes the romanized form of text as an interface for LLMs, hypothesizing that its frequent informal use and shared tokens with English enhance cross-lingual alignment. Our approach involves the continual pretraining of an English LLM like Llama 2 (Touvron et al., 2023) on romanized text of non-English, non-Roman script languages, followed by instruction tuning on romanized data. The results indicate that romanized text not only reduces token fertility by 2x-4x but also matches or outperforms native script representation across various NLU, NLG, and MT tasks. Moreover, the embeddings computed on romanized text exhibit closer alignment with their English translations than those from the native script. Our approach presents a promising direction for leveraging the power of English LLMs in languages traditionally underrepresented in NLP. Our code is available on https://github.com/AI4Bharat/romansetu.
[…]
[backgroundcolor=gray!10, linecolor=gray!70, linewidth=1pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, font=, ]promptbox
@@
Jaavid Aktar Husain1,2†§  Raj Dabre1,5,8,9§  Aswanth Kumar3  Jay Gala4‡ Thanmay Jayakumar1Ratish Puduppully6Anoop Kunchukuttan1,7,8§ 1Nilekani Centre at AI4Bharat  2IIIT D&M Kancheepuram  3Flipkart 4Mohamed bin Zayed University of Artificial Intelligence 5National Institute of Information and Communications Technology, Kyoto, Japan 6Institute for Infocomm Research (I2R), ASTAR, Singapore 7Microsoft, India  8Indian Institute of Technology Madras, India 9Indian Institute of Technology Bombay, India
@@ 1 Introduction
Large Language Models (LLMs) demonstrate remarkable proficiency across a broad spectrum of natural language processing (NLP) tasks, as evidenced by various studies (Liu et al., 2023; Chung et al., 2022; Chowdhery et al., 2023; Wei et al., 2024; Goyal et al., 2022b). They excel not only in tasks for which they were explicitly trained but also in those for which they were not trained. This achievement is mainly due to the availability of corpora (Wenzek et al., 2020; Abadji et al., 2021; Suárez et al., 2019) as well as the advancements in LLMs that leverage these datasets for pretraining (Touvron et al., 2023; Workshop et al., 2022; Chowdhery et al., 2023). Despite their proficiency in English, these models typically demonstrate reduced effectiveness when applied to non-English languages, highlighting a significant challenge in extending their benefits to non-English languages.
@@
Figure 1: Performance of the BaseLLM (Llama 2 Base), our continually pretrained model (CPT), and our instruction finetuned model (IFT), in both native (N) and romanized (R) script settings. The average scores for different tasks (left) and languages (right) are compared in each radar chart. For CPT, we show 3-shot results where available, or else 1-shot results. For IFT, we use zero-shot results.
@@
The English-heavy LLMs (Touvron et al., 2023; Jiang et al., 2023; Zhang et al., 2022) still have some representation coverage from other languages due to data leakage while creating the pre-training dataset, particularly for languages that use the same script as English i.e., the Roman script. This script sharing enables cross-lingual transfer and bestows some of these LLM capabilities to these languages. For languages using non-Latin scripts, the data representation is very limited to non-existent. Tokenization of text in the languages exhibits high fertility (Ács, 2020) and byte-level representation (Artetxe et al., 2020) due to inadequate representation of these languages in the tokenizer vocabulary. Hence, these LLMs perform poorly on most of these languages, and inefficient tokenization also leads to high inference latency as well as the inability to process long sequences. This disparity in performance raises a critical question: how can we extend the capabilities of LLMs to the languages written in non-Latin scripts?
@@
A widely explored solution is the extension of the tokenizer vocabulary to incorporate new languages and continual pre-training on native language data (Cui et al., 2023; Nguyen et al., 2023; Minixhofer et al., 2022). This approach is computationally demanding since the models need to be continually pre-trained long enough to effectively learn the new embeddings and align the representations of English and the new language. Furthermore, this approach requires the availability of large volumes of text corpora.
@@
In this work, we explore an alternative approach for more efficient knowledge transfer. Instead of utilizing the native script, we use the romanized form of text as the interface to the LLMs in our approach RomanSetu. [^1] The word setu in Sanskrit means bridge, referring to the Roman script serving as a bridge between English and other languages. The adoption of romanized representation is justified for several reasons. In many languages, romanized text has been frequently used in informal settings and on social media in recent history. This usage creates the potential for the inclusion of some romanized data representation in the pre-training corpus. Additionally, code-mixing with English is a common occurrence and the romanized form shares tokens with English. This leads us to hypothesize that the romanized form is better aligned with English than the native script, thereby facilitating a more effective transfer from English.
@@
In our approach, we first continually pre-train an English LLM like Llama 2 (Touvron et al., 2023) on romanized text for a language generated via transliteration. Subsequently, we perform instruction tuning on this continually pretrained model. We experiment with 5 languages from two language families and evaluate our approach on several NLU, NLG and MT benchmarks. Our experiments and analysis indicate that:
@@
  • Results across several NLU, NLG, and MT tasks show that multilingual instruction finetuning on romanized data produces competitive or superior results than instruction finetuning on native script data, highlighting efficiency and better cross-lingual transfer. Figure 1 summarizes these results of various models on romanized and native inputs.
@@
  • Romanized representation can enable cross-lingual transfer in decoder-only English-heavy LLMs. Previous work has primarily focused on the use of romanization in multilingual, encoder-only models.
@@ 2 Related Work
In NMT, Amrhein and Sennrich (2020) show that transliteration shows improvements for low-resource languages with different scripts by transferring from related high-resource languages that use different scripts. Goyal et al. (2020) show that transliteration helps even when only a contact relationship exists between the languages involved. Song et al. (2020) show that transliteration during the pretraining stage for NMT also helps cross-lingual transfer.
Transliteration has also been used for cross-lingual transfer in the context of pretrained language models. Some works (Khemchandani et al., 2021; Dhamecha et al., 2021; Moosa et al., 2023; Purkayastha et al., 2023) employ transliteration to a common script during the pretraining phase to enable cross-lingual transfer. Other studies (Dabre et al., 2022; Muller et al., 2021; Chau and Smith, 2021) adopt transliteration during the fine-tuning. Transliteration could be done to a common non-Latin script (Khemchandani et al., 2021; Dhamecha et al., 2021; Dabre et al., 2022; Doddapaneni et al., 2023) or to the Latin script Muller et al. (2021); Moosa et al. (2023); Purkayastha et al. (2023).
@@
  • Previous work explored cross-lingual transfer using transliteration with multilingual language models. We focus our attention on English LLMs and try to achieve cross-lingual transfer via romanization using English LLMs. This is a challenging scenario since the language can be unrelated to English, and very little (if any) native or romanized data in the language under study might be seen during pre-training of the English LLM. At the same time, this is a very practical need since most best performing LLMs are English-heavy Touvron et al. (2023); Jiang et al. (2023) and hence cross-lingual transfer via romanization is an important direction to explore.
@@
To the best of our knowledge, no prior research has investigated leveraging of romanization for mostly English LLMs in the context of cross-lingual transfer to non-English languages.
@@ 3 Utilizing Romanized Data to make LLMs Multilingual
Our proposed approach aims to enhance the capabilities of mostly English LLMs for non-English languages by leveraging romanized data. We first continually pretrain the LLM with both romanized and English data to create a base LM that is romanization-aware. Subsequently, we extend the process by instruction tuning the continually pretrained model. The framework of our approach is illustrated in Figure 2, encompassing the stages of data romanization, pretraining, and instruction tuning.
@@ 3.1 Romanization Scheme
A variety of romanization schemes are available, each driven by multiple considerations. One key factor is the resemblance of the romanized representation to the way people typically write romanized text. This is particularly advantageous if the pretraining data includes romanized text, as it might aid in aligning with English. Another important aspect is the fertility achieved by the romanization scheme when considering the original LLM’s tokenizer. A third consideration is whether the transliteration scheme is lossy or lossless. A lossless scheme is preferable when the objective is to convert the output back to the native script. Typically, deterministic transliteration schemes are lossless, whereas natural transliteration schemes tend to be lossy.
In this work, we have focused on Indic languages. Hence, we evaluated two romanization schemes for Indic languages: (a) the extended ITRANS scheme from the IndicNLP library Kunchukuttan (2020), which defines a fixed, reversible mapping between Indic scripts and Roman characters, and (b) the IndicXlit scheme Madhani et al. (2023), which generates romanizations as commonly used by Hindi speakers in informal contexts and is learned from parallel transliteration corpora. These mappings are inherently lossy. The IndicXlit romanization demonstrates lower fertility compared to ITRANS romanization (See Table 1). Preliminary prompting experiments with the Llama 2 model also indicated that IndicXlit outperforms ITRANS in romanized Hindi to English translation. Consequently, we selected IndicXlit as our romanization scheme for this project. However, it is important to note that the transliterations are not reversible. With continued pretraining, ITRANS transliterations might eventually achieve similar task performance to IndicXlit while ensuring script reversibility. This direction might also be relevant for other languages which do not have a transliteration model available, while romanization mappings are available for almost all scripts. We leave this exploration for future work.
@@ 3.2 Continual Pretraining
While the base model is trained on roman script, most of the data it has been exposed to is in English. Hence, the base model cannot be used as-is for processing romanized inputs, particularly for generation into non-English languages. Hence, the base model needs to be updated. To make the base model romanization-aware, we continue to pretrain the base LLM with romanized document-level data. To prevent any catastrophic forgetting of English capabilities, we also incorporate an equal amount of English data into the pretraining mix. We perform full-finetuning of the models. We do not expand or change the model vocabulary, hence the romanized inputs can take advantage of the token embeddings from the base model while adapting to the new languages.
@@ 3.3 Instruction Finetuning (IFT)
Non-English languages possess very limited finetuning data. A common method to overcome this limitation is to translate English IFT datasets into the languages. We could choose to translate both input and outputs Wei et al. (2023), or translate only the inputs and ask a powerful, proprietary LLM like ChatGPT to generate translations in the native language Li et al. (2023). We chose the former approach since the quality of LLM responses (even for the best commercial models) might not be of good quality Ahuja et al. (2023), while high-quality open-source translation models Costa-jussà et al. (2024); Gala et al. (2023) are available making translation of both inputs and outputs feasible at scale for many languages. The translated IFT datasets are then romanized using IndicXlit.
@@ 4.1 Languages
We experimented with the following Indic languages: Hindi, Marathi, Gujarati, Tamil and Malayalam. These languages span 2 language families and 4 different scripts and are mid to low-resource languages. The languages considered in this study belong to the Indo-Aryan branch of Indo-European (Hindi, Marathi, Gujarati) and Dravidian (Tamil, Malayalam) families. In these languages, the typical word order is SOV. These languages have an inflectional morphology, with the Indo-Aryan leaning somewhat towards fusional, while Dravidian languages tend to have an agglutinative morphology. It is important to note that the script systems used for these languages are abugida, effectively faithfully capturing vowels and consonants. We see the choice of Indian languages as a strength since it spans multiple language families, multiple scripts, and diverse linguistic characteristics and resource levels. Past work has also used Indic languages as a case study for romanized representation Moosa et al. (2023); Dhamecha et al. (2021).
@@ 4.2 Datasets
Instruction Fine-tuning: These languages have very little native instruction tuning data for diverse tasks. Following Wei et al. (2023), we rely on translating high-quality English-supervised instruction-tuning datasets into the languages under consideration. We use IndicTrans2 (Gala et al., 2023), the state-of-the-art open-source MT model for Indian languages compared to commercial offerings for translation. We sampled examples from various English instruction tuning datasets to ensure a diverse mix of tasks from the Flan collection (Longpre et al., 2023) (65k) and Dolly (Conover et al., 2023) (15k). These are translated into all the languages. The instruction tuning dataset includes a translation subset as well to learn the translation task as well as to drive alignment between English and other languages. We sample 20k high-quality manually translated examples in each direction from the BPCC-Human subset (Gala et al., 2023). Thus, the final IFT dataset has 120k examples per language. The instruction tuning datasets are further romanized using IndicXlit.
Evaluation Data: We evaluate our models on a variety of NLU and NLG tasks. These include native language test sets from the IndicXTREME (Doddapaneni et al., 2023) and Indic NLG Suite (Kumar et al., 2022) benchmarks. For the translation task, we use the FLORES-200 devtest (Goyal et al., 2022a; Costa-jussà et al., 2024) for evaluation. Following recent work (Lai et al., 2023; Üstün et al., 2024; OpenAI et al., 2023), we evaluate the knowledge and reasoning capabilities of the model on the machine-translated versions of the benchmarks such as MMLU (Hendrycks et al., 2021a), ARC (Clark et al., 2018a), BoolQ (Clark et al., 2019a) and CommonSenseQA (Talmor et al., 2019). Note that as the CommonSenseQA benchmark comprises of a blind test set, we report the evaluations on the dev set. The translations were obtained using IndicTrans2 and the romanized inputs were created using IndicXlit.
@@ 4.3 Training and Finetuning Details
The models are fully finetuned during continual pre-training as well as supervised fine-tuning. Both the continued pretraining and instruction fine-tuning are multilingual i.e. a single model is trained jointly with data for all languages (including English to preserve English performance). Note that no vocabulary expansion is done. For pre-training, the loss is computed for all the tokens in the sequence. During instruction fine-tuning, the loss was computed only for the output tokens. Detailed information about the hyperparameters used for both continual pretraining and instruction fine-tuning can be found in Appendix A. We adapt open-instruct [^2] https://github.com/allenai/open-instruct/tree/main (Wang et al., 2023) for our continual pretraining and fine-tuning experiments.